OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (9, 7) and (4, –8). y − = (x − 9)

OpenStudy (anonymous):

@Directrix

OpenStudy (diamondboy):

ok do u know d point slope formula?

OpenStudy (anonymous):

no

OpenStudy (diamondboy):

\[(y-y1)=m(x-x1)\]

OpenStudy (diamondboy):

ok

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 9}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 4}}\quad ,&{\color{blue}{ -8}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

i got 1/5

OpenStudy (diamondboy):

your y1 is 7 and your x1 is already given as 9

OpenStudy (diamondboy):

all u need is look for d slope

OpenStudy (jdoe0001):

got slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can i has it pleaseeeeeee

OpenStudy (jdoe0001):

well.. that's the idea of the exercise, that you find it

OpenStudy (jdoe0001):

so \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 9}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 4}}\quad ,&{\color{blue}{ -8}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\) so..... what would that give you for "m" or the slope?

OpenStudy (anonymous):

2/15

OpenStudy (anonymous):

so simplify it for answer

OpenStudy (diamondboy):

ok what is ur y2 and y1?

OpenStudy (anonymous):

-8 is my y2 my y1 is 7

OpenStudy (diamondboy):

ok subtract dem

OpenStudy (anonymous):

i did i got -15

OpenStudy (diamondboy):

ok

OpenStudy (diamondboy):

what is ux2 and x1

OpenStudy (diamondboy):

sorry r x2

OpenStudy (diamondboy):

whatever u get subtract dem

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-17

OpenStudy (diamondboy):

now divide -15/-17

OpenStudy (anonymous):

i cant get the answer